### v1

# Holistic Tokenizer for Autoregressive Image Generation

Abstract
The vanilla autoregressive image generation model generates visual tokens in a step-by-step fashion, which limits the ability to capture holistic relationships among token sequences. Moreover, most visual tokenizers map local image patches into latent tokens, leading to limited global information. To address this, we introduce Hita, a novel image tokenizer for autoregressive (AR) image generation. It introduces a holistic-to-local tokenization scheme with learnable holistic queries and local patch tokens. Besides, Hita incorporates two key strategies for improved alignment with the AR generation process: 1) it arranges a sequential structure with holistic tokens at the beginning followed by patch-level tokens while using causal attention to maintain awareness of previous tokens; and 2) before feeding the de-quantized tokens into the decoder, Hita adopts a lightweight fusion module to control information flow to prioritize holistic tokens. Extensive experiments show that Hita accelerates the training speed of AR generators and outperforms those trained with vanilla tokenizers, achieving 2.59 FID and 281.9 IS on the ImageNet benchmark. A detailed analysis of the holistic representation highlights its ability to capture global image properties such as textures, materials, and shapes. Additionally, Hita also demonstrates effectiveness in zero-shot style transfer and image in-painting. The code is available at https://github.com/CVMI-Lab/Hita.
1 Introduction
Within an autoregressive (AR) generation paradigm, the field of large language models (LLMs) [51, 9, 37, 35, 4, 61, 52] has witnessed significant progress in recent years. In particular, using AR generation along with transformer backbones, GPT-style models have exhibited impressive performance [31, 13, 1], incredible scalability [22, 16], and versatile flexibility [35, 4] across various language tasks.
Figure 1: The concept of holistic tokenizer. A set of learnable queries that capture global properties, such as color, texture, material, etc, from pixels, with semantic-level feature injected, is utilized to reconstruct the image along with image patches.
Figure 2: Holistic tokens capture the global style and content information, as demonstrated by their ability to enable (a) style transfer by replacing the holistic tokens of input images with those of reference images, and (b) zero-shot image inpainting using pre-trained AR generation models . Additionally, the introduction of these new latent tokens accelerates the training speed of the GPT-style AR image generation model by a factor of two (c).
The success of autogressive models in language modeling has driven substantial interest in their application to image generation, giving rise to representative works such as DALL-E [38], Parti [55, 56], VAR [47], and LlamaGen [46]. In AR-based image generation, VQVAE [12] is employed to encode images into discrete tokens. These tokens are then raster-scanned into a 1D sequence to train an AR transformer model like Llama [48], which predicts the next token in the sequence. During inference stage, the token sequence generated by the AR model is decoded by the VQVAE’s [12] decoder to produce the synthesized images.
While significant advancements have been made, most research efforts have concentrated on refining and scaling model architectures or optimizing learning objectives for either the VQVAE stage [46, 55, 56, 58] or the AR transformer stage [46, 38]. However, in terms of latent space design, existing approaches predominantly rely on patch-level image representations at either single scales [56, 46] or multiple scales [47], which limits the ability of latent tokens to capture global and semantic-level information. Although these spatially structured tokens are compatible with diffusion models [44, 10, 19, 33, 41, 32], which predict all tokens in parallel, they introduce challenges for AR-based models. Autoregressive models typically employ causal attention to predict tokens sequentially [6, 49, 47, 46]. The absence of holistic information makes it difficult for these models to maintain global coherence, leading to limitations in their capabilities and increasing learning difficulties. As shown in Fig. 2b (LlamaGen[46]) with image inpainting, when half of an image is input into the AR model as a prefix prompt, requiring it to generate the remaining part, the completed images do not align well with the original content. This challenge has also been highlighted in works such as [11, 27, 53], where bi-directional attention mechanisms are introduced to mitigate the issue, albeit with non-trivial modifications to the AR generation process.
To address this, we introduce Hita, a novel global-to-local image tokenizer for AR image generation. First, we propose to utilize a set of learnable queries to capture the holistic information from the patch-level embeddings of an image. Meanwhile, a pre-trained foundation model, such as DINOv2 [7], is adopted to inject semantics into the learned queries to enhance their representation. Then, we utilize 2 separate codebooks to quantize the learned queries and patch-level embeddings, respectively, resulting in holistic tokens and patch-level tokens. Next, we fuse the de-quantized holistic and patch-level tokens together with a novel token fusion module to reconstruct the image. Such design can better prioritize the holistic token and improve both image reconstruction and generation quality.
Once Hita is trained, a vanilla autoregressive (AR) transformer model, e.g. Llama [48], can be seamlessly incorporated to generate both holistic and image patch tokens sequentially, which can be decoded into image facilitated by Hita. During the AR generation, the AR generator first generates holistic tokens, which encapsulate the global image information and serve as a prefix prompt. These holistic tokens then guide the subsequent generation of local image patch tokens, alleviating difficulties and maintaining global coherence to improve generation quality.
Our in-depth ablation experiments confirm that the learned holistic tokens effectively capture global image features, such as shape, color, texture, etc, from pixels (see Fig. 2a). This has been demonstrated through effortlessly enabling style transfer by combining holistic tokens from reference images with patch-level tokens from the source image. Moreover, without task-specific training, the model can be seamlessly applied to image inpainting by encoding the given incomplete image into latent tokens, which serve as prefix sequence prompts to the AR model, requiring it to generate visual tokens for completing the image (see Fig. 2b), demonstrating the capability of holistic tokens in ensuring global consistent synthesis.
To further validate its effectiveness, we benchmark Hita for AR image generation using the leading Llama-based [48] AR transformer model. Extensive experiments show that Hita accelerates the convergence speed during training and significantly improves synthesis quality (see Fig. 2c). Specifically, the time required to reach an FID of 4.22 is reduced by a factor of 2.1. Furthermore, our final AR model with 2B parameters achieves an FID of 2.59 on the ImageNet 256×256 class-conditional image generation benchmark. This outperforms the well-known AR generation model LlamaGen [46] with 3B parameters. It also surpass the popular diffusion model LDM-4[41] by 0.9 FID and 27.1 IS, respectively.
To summarize, our contribution can be listed as follow:
  • We propose a novel global-to-local image tokenizer for AR image generation, with which an AR generation model can be integrated without any modification.
  • The proposed image tokenizer emerges with new characteristics, such as zero-shot style stransfer and zero-shot image in-painting.
  • Extensive experiments conducted on class-conditional image generation task demonstrates it can accelerate the converge speed during training and significantly improve the synthesis quality.
2 Related Work
Image Tokenization using Autoencoders. Image tokenization aims to transform images into compact latent tokens, reducing computation and redundancy for generative models [57, 29, 47, 46]. VQVAEs [50, 40] integrate vector quantization into the VAE framework [23] to transform continuous visual signals into discrete tokens, thus enabling autoregressive modeling like LLMs. Based on VQVAEs [50, 40], VQGAN [12] improves compression and reconstruction by adding an adversarial loss, while ViT-VQGAN [54] combines transformers with VQGAN. Further advancements, like RQ-VAE [26] and MoVQ [62], explore multiple vector quantization stages for better latent embeddings. Recently, MAGVIT-v2 [29] and FSQ [30] introduced quantization methods that do not rely on a lookup codebook. VQGAN-LC [63] clusters features from a pre-trained CLIP [36] model, expanding the token set to over 10,000 with 99% utilization. In contrast to methods that focus on learning patch – level latent tokens– often lacking global context and coherency, which can limit the effectiveness of AR models– our work emphasizes the importance of holistic representation in the latent tokens.
TiTok [59] and VAR [47] are related efforts but differ from Hita in several aspects. TiTok aims to create compact 1D tokens with queries to reduce redundancy, while Hita uses queries to capture holistic properties for 1D holistic tokens and fuses them with 2D image patch tokens to reconstruct images. TiTok refines queries for 1D tokens tailored for reconstruction objectives, whereas Hita extracts information from pre-trained foundation model and patch features to capture global information. TiTok requires a complex training strategy, while Hita enables efficient single-stage training. Hita outperforms TiTok in AR generation and offers new features like zero-shot style transfer and inpainting. VAR [47] predicts global and local tokens indiscriminately, while Hita separates holistic and patch-level tokens, enriching holistic tokens with pre-trained foundation to form global representations and uses causal attention to align with the causal nature of AR models, which VAR lacks. Removing initial coarse-scale tokens seldom affects VAR’s reconstruction, indicating they are not crucial.
Autoregressive Image Generation. Recent advancements in GPT-style transformers have sparked significant interest in AR generation, leading to various representative approaches [6, 49, 46, 26, 47, 28]. Like GPT [34], these models predict sequential visual tokens autoregressively. Early AR models [6, 49] generated sequences directly in pixel space, while current models [26, 54, 47, 46] use tokenizers [50, 40, 54, 12] to compress images into discrete tokens, followed by causal transformers for next-token prediction and image reconstruction through the tokenizer’s decoder. Most tokenization techniques generate discrete visual tokens in a patch-wise fashion, assuming a 2D grid structure for image signals. While logical for spatially structured inputs, this patch-based approach limits the tokenizer’s ability to capture holistic information. This limitation is amplified in AR models, where causal attention’s step-by-step prediction struggles to capture long-range dependencies. Studies [3, 11, 27] show that causal attention in image token generation underperforms compared to bidirectional attention.
To address this, recent works [47, 27, 53] have introduced bidirectional attention to AR models with promising results. VAR [47] introduced next-scale prediction, enabling simultaneous token generation and allowing bidirectional attention within each scale. MAR [27] incorporated a BERT-style framework to integrate bidirectional attention, while Show-o [53] combined unidirectional and bidirectional attention in a single transformer for both comprehension and generation. However, these approaches complicate the design of a universal transformer to unify multi-modal understanding and generation while maintaining the next-token prediction paradigm. In this work, we introduce a new global-to-local tokenization method and investigate how far we can advance AR generation while preserving the inherent design of language models.
3 Method
In this section, we first introduce the preliminary related to image tokenizer and AR generation models. Then, we present the design of our holistic tokenizer step by step. Next, their characteristics are also discussed.
Figure 3: The flowchart of the proposed holistic image tokenizer, which includes: Holistic feature extraction part incorporates learnable queries to obtain holistic and patch-level features from pixels. Separate vector quantization part process holistic and patch tokens by separate codebooks. Token fusion and decoding part leverages a lightweight token fusion module to fuse holistic tokens into patch-level tokens to reconstruct image. Here we omit de-quantization between quantizer and token fusion module for simplicity.
3.1 Background
Quantized Image Tokenizer. To apply autoregressive modeling to visual generation, existing methods [55, 56, 46, 47] necessitate an image tokenizer to convert a 2D image into a 1D token sequences. Quantized autoencoders, such as VQVAEs [50, 40, 12, 54, 63, 47, 46], are widely used. An image tokenizer generally consists of an encoder $\mathcal{E}(\cdot)$, a quantizer $\mathcal{VQ}(\cdot)$, and a decoder $\mathcal{D}(\cdot)$. Given an input image $\text{I}\in\mathbb{R}^{H\times W\times 3}$, the encoder $\mathcal{E}(\cdot)$ first projects image pixels to the feature map $Z_{2D}\in\mathbb{R}^{\frac{H}{f}\times\frac{W}{f}\times\text{D}}$ with spatial down-sampling factor $f$. Then, $Z_{2D}$ is fed into the quantizer $\mathcal{VQ}(\cdot)$ that typically includes a learnable codebook $\mathbb{C}\in\mathbb{R}^{N\times D}$ with $N$ vectors. Each feature vector ${z_{i}}\in\mathbb{R}^{D}$ is mapped into its nearest vector ${c_{i}}\in\mathbb{R}^{D}$ in the codebook $\mathbb{C}$. This process can be formulated as Eq. (1).
$\begin{split}{Z_{2D}}&=\mathcal{E}(\text{I}),\\ \mathcal{VQ}({z_{i}})={c_{i}},\quad\text{where}\quad{i}&=\mathop{\arg\min}\limits_{{j}\in\{1,2,...,N\}}\|{z_{i}}-{c_{j}}\|_{2}.\end{split}$ (1)
where $H$ and $W$ denote the input image’s height and width respectively. $D$ depicts the latent feature dimension. Once discrete tokens are acquired, they can be de-quantized into corresponding code and converted back to image pixels by the decoder $\mathcal{D}(\cdot)$, as depicted in Eq. (2).
$\mathbf{\hat{I}}=\mathcal{D}(\mathcal{VQ}({Z_{2D}})).$ (2)
For the optimization of codebook, the training objective is $\mathbf{\mathcal{L}_{vq}}=\sum{\|\mathbf{sg}(z_{i})-{c_{i}}\|^{2}_{2}+{\beta}\cdot\|\mathbf{sg}({c_{i}})-{z_{i}}\|^{2}_{2}}$, where $\mathbf{sg}(\cdot)$ is a stop-gradient function [2, 50]. The second term is a commitment loss with loss weight $\beta$ used to align extracted features with codebook vectors. For image reconstruction optimization, the loss function is $\mathcal{L}_{AE}=\mathcal{L}_{2}(\text{I},\hat{\text{I}})+\mathcal{L}_{P}(\text{I},\hat{\text{I}})+\lambda_{G}\cdot\mathcal{L}_{G}(\hat{\text{I}})$, where $\mathcal{L}_{2}$ is a pixel-wise reconstruction loss, $\mathcal{L}_{P}$ is perceptual loss from LPIPS [60], $\mathcal{L}_{G}$ is adversarial loss from PatchGAN [20] with weight $\lambda_{G}$.
Autoregressive Image Generation. When used for visual generation, the 2D quantized map from a tokenizer is converted into a 1D sequence in a pre-defined order, e.g. raster-scan order. Then, an AR generation model predicts the sequence of image tokens autoregressively. To achieve this, AR models typically incorporate causal attention with prefix token sequence to predict the next, which focuses on learning the dependencies between the prefix discrete image tokens and the current one. Considering a sequence of discrete tokens ${x_{1:N}}$, an AR model is trained on these tokens by maximizing the cross-entropy depicted in Eq. (3), where ${c}$ is class label or text embedding, ${x_{j<i}}$ serves as the prefix sequence of tokens for the current token ${x_{i}}$.
$\mathop{\max}\mathcal{L}^{\text{CE}}_{\theta}=\sum\limits_{i=1}^{N}\text{log}P_{\theta}({x_{i}}|{x_{j<i}},c)$ (3)
During inference, the AR model initially predicts visual tokens by sampling tokens from the learned distribution autoregressively. Subsequently, the generated visual tokens are converted back into images through the decoder $\mathcal{D}(\cdot)$.
3.2 Holistic Tokenization for AR Generation
As shown in Fig. 3, Hita incorporates the following key components. First, to capture holistic property, we introduce learnable queries that interact with image patch embeddings via attention modules. A pre-trained foundation model (e.g., DINOv2 [7]) injects semantic-level features. To align Hita’s latent space with the causal nature of AR models, we use a causal transformer to integrate holistic and patch features before quantization, embedding the sequential structure into the tokenization process to learn prefix-aware latent tokens. Details are depicted in Sec. 3.2.1. Then, the holistic and patch-level features are quantized using separate codebooks (see Sec. 3.2.2). Finally, in the token fusion and decoding procedure after de-quantization, Hita first integrates a fusion module with casual transformers to process these tokens in order and prioritize holistic tokens. Then they are fed into the decoder for image reconstruction (see Sec. 3.2.3).
3.2.1 Holistic Feature Extraction
We build Hita based on VQGAN [12]. Specifically, we adopt its encoder $\mathcal{E}(\cdot)$ to encode input image $I$ into local image patch embeddings, and further incorporate a set of learnable queries ${Q}$ to gather holistic information from the whole input. To enhance the holistic representation, a pre-trained foundation model $\mathcal{H}(\cdot)$ like DINOv2 [7] is leveraged to inject rich semantic-level features. To achieve this, we first concatenate the learnable queries and the flattened patch-level embeddings as well as foundation model features into sequence and feed them into a simple transformer $\mathcal{E}_{\text{{trans}}}(\cdot)$ is utilized, where holistic queries ${Q}$ interact with all patch-level image embeddings and the injected semantic-level features. Considering the AR model models the probability of observing the current token based on its predecessors, a well-aligned latent space with the causal nature is encouraged. Therefore, the extracted holistic feature and patch-level feature are concatenated and reorganized in an ordered sequence that holistic queries come first followed by the patch-level image patch feature obtained in raster-scan order. This sequence is then fed into a transformer with global features serving as prefix prior. The whole process is formulated as Eq. (4).
$\begin{split}{Q},\ {Z}&=\mathcal{E}_{\text{trans}}({Q}\oplus\mathcal{E}({I})\oplus\mathcal{H}({I}))\\ {\overline{Q}},\ {\overline{Z}}&=\mathcal{E}_{\text{casual}}({Q}\oplus{Z})\end{split}$ (4)
where ${Z}\in\mathbb{R}^{\frac{H}{f}\times\frac{W}{f}\times{C}}$ refers to the path-level embeddings with a spatial downsample ratio $f$. ${Q}\in\mathbb{R}^{M\times C}$ denotes learnable queries. $\oplus$ is the concatenation operation. Moreover, the pre-trained feature is discarded after $\mathcal{E}_{\text{trans}}$.
Figure 4: The illustration of token fusion and decoding procedure
3.2.2 Separate Vector Quantization
Once both learned queries $\overline{Q}$ and patch-level images $\overline{Z}$ are produced, they are quantized with holistic quantizer $\mathcal{Q}_{H}(\cdot)$ and patch quantizer $\mathcal{Q}_{P}(\cdot)$, respectively. Then, the holistic tokens $\hat{Q}$ and patch-level $\hat{Z}$ tokens could be obtained. Considering the codebook significantly impacts the performance of the image tokenizer. Following [54, 46], we apply ${\ell}_{2}$-normalization to the codebook vectors, opting for a lower vector dimension and a larger codebook size to improve the reconstruction quality and codebook utilization.
3.2.3 Token Fusion and Decoding Procedure
After quantization, the holistic tokens and patch-level tokens could be fed into a vanilla transformer to perform token fusion and then fed into the decoder to reconstruct the image. However, such a design incurs codebook collapse in the holistic quantizer (3th row in Table. 3). This occurs because patch-level tokens directly impact the reconstruction of relevant image patches through the skip connection in the transformer, which makes it bypass holistic tokens (see Fig. 1a in Appendix). As a result, the token fusion module learned a trivial solution, which overlooks the holistic tokens and leads to holistic codebook collapse.
[…]
3.2.4 Training objective
For tokenizer optimization, we adopt the training losses used in VQGAN [12, 46], except for the slight difference in the vector quantization loss $\mathcal{L}_{vq}$. Given that we use two separate codebooks to quantize the holistic and local patch features independently, the vector quantization loss $\mathcal{L}_{vq}$ can be formulated as $\mathcal{L}_{vq}=\mathcal{L}_{vq}(\overline{Q})+\mathcal{L}_{vq}(\overline{Z_{p}})$. Consequently, the total losses for guiding the training of our image tokenizer can be represented as $\mathcal{L}=\alpha\cdot\mathcal{L}_{vq}+\lambda\cdot\mathcal{L}_{AE}$. In this case, we set both $\alpha$ and $\lambda$ to 1.
3.3 Autoregresive Image Generation
Once the tokenizer is trained, a standard autoregressive (AR) transformer model, e.g. Llama [48], can be seamlessly integrated without requiring any revision to generate these latent tokens sequentially. Starting from the class/text embedding $c$, the AR model first generates holistic tokens as prefix prompts, then guides the subsequent generation of patch token sequence in the way of next-token prediction. Once all the visual tokens are generated, they will be tackled by the token fusion and decoding procedure to be converted into pixels. Additionally, in AR models, each visual token is added with positional embedding. Since holistic tokens form a 1D sequence, we add learnable positional embeddings for them. Meanwhile, image patch tokens continue to use 2D RoPE [45] for their positional embeddings, which remains the same as LlamaGen [46]
3.4 Discussions
What do holistic tokens capture? To understand what has been learned by the holistic tokens, we perform an experiment on the learned holistic tokens. Specifically, two images are first encoded and quantized into holistic tokens and image patch tokens, respectively. Then, the holistic tokens from one image are concatenated with the patch tokens from the other. These tokens are fed through the de-quantization process and the decoder to synthesize a new image. This entire procedure is conducted using a pre-trained holistic tokenizer, without any additional model training or fine-tuning. As shown in Fig. 2a, the holistic tokens can capture global information from an image, such as texture, color, materials, shapes, and other features from the pixels. For instance, in the case of the lizard image, Hita successfully transfers the material properties of the ship from the reference image in thumbnail to the input image, creating a realistic mechanical lizard. More samples are shown in Appendix.
Does holistic tokenizer enhance semantic coherence? To verify whether the holistic features can enhance the global coherence of the generated content, we perform a zero-shot image in-painting evaluation. In this in-painting task, the model is forced to complete the lower half of an image given only the upper half. This setup prevents local information leakage during in-painting. We first tokenize the images into holistic and patch tokens, then feed them to the AR model to generate additional tokens, conditioned on class embeddings in AR model. Hita-XXL is used without modifications to the model architecture or parameter tuning.
As shown in Fig. 2b, Hita effectively completes the lower part of the image based on the upper part, producing images that maintain strong semantic consistency. In contrast, outputs from LlamaGen [46] lack this consistency. For example, in the case of the fish image, Hita correctly generates a complete fish, while LlamaGen [46] mistakenly generates a “fish-bird”. These results confirm that our holistic tokenizer is semantic-aware and has a stronger ability to generate semantically and holistically coherent examples. Furthermore, the ablation study of linear probing on holistic tokens is conducted in Sec. 4.
4 Experiments
[…]
GAN GigaGan [21] 569M 3.45 225.5 0.84 0.61
StyleGan-XL [43] 166M 2.30 265.1 0.78 0.53
Diff. LDM-4 [41] 400M 3.60 247.7 0.87 0.48
DiT-XL/2 [32] 675M 2.27 278.2 0.83 0.57
Mask. MaskGIT [5] 227M 256 6.18 182.1 0.80 0.51
MaskGIT-re [5] 4.02 355.6
AR TiTok-Base [59]  [^1] M 256 50 9.19 170.4 0.85 0.33
TiTok-Large [59] 343M 5.66 206.4 0.84 0.45
GPT2-re [12] 1.4B 256 5.20 280.3
VIM-L-re [54] 1.7B 1024 3.04 227.4
RQTran.-re [26] 3.8B 68 3.80 323.7
LlamaGen-B [46]  [^1] M 576 50 8.31 154.7 0.84 0.38
LlamaGen-L [46] 343M 4.24 206.7 0.83 0.49
LlamaGen-XL [46] 775M 3.24 245.7 0.83 0.53
LlamaGen-XXL [46] 1.4B 2.89 236.2 0.81 0.56
LlamaGen-3B [46] 3B 2.61 251.9 0.80 0.56
Hita-B  [^1] M 569 50 5.85 212.3 0.84 0.41
Hita-L 343M 3.75 262.1 0.85 0.48
Hita-XL 775M 2.98 253.4 0.84 0.54
Hita-XXL 1.4B 2.70 274.8 0.84 0.55
Hita-2B 2B 2.59 281.9 0.84 0.56
Titok-B [59]  [^1] M 256 300 6.91 173.3 0.83 0.42
LlamaGen-B [46] 576 6.09 182.5 0.85 0.42
Hita-B 569 4.33 238.9 0.85 0.48
Titok-L [59] 343M 256 300 4.00 242.1 0.84 0.50
LlamaGen-L [46] 576 3.07 256.1 0.83 0.52
Hita-L 569 2.86 267.3 0.84 0.54
Table 1: Class-conditional image generation quality on Imagenet [8] benchmark. indicates the approach is implemented by us ‘-re’ indicates using rejection sampling.
4.1 Setup
Image Tokenizer.
Following [46, 59], we set the codebook vector dimension of the patch-level image token and holistic token to 8 and 12, respectively. As discussed in [46], this design can achieve a better reconstruction and efficient codebook usage, Besides, both codebooks have a size of $16,384$. The number of queries for holistic feature capture is fixed at 128. Moreover, to preserve the tokenizer’s simplicity, the depth of each transformer is set to 3, and the holistic token selection length $k$ is set to 4, by default.
Hita is optimized on the ImageNet [8] training set and evaluated on the validation set. To ensure a fair comparison with LlamaGen [46], we train the tokenizer on $336\times 336$ images, which yields a comparable number of discrete tokens to it. We also adopt the training settings for tokenizer training from LlamaGen. Additionally, during evaluation, the test images are resized to $256\times 256$, consistent with the evaluation procedure in LlamaGen.
Class-conditional Autoregressive Image Generation.
Following the AR generation procedure in LlamaGen [46], the AR models first generate images of the same size as the image tokenizer and then resize them to $256\times 256$ for evaluation. Unless otherwise specified, all models are trained with the same settings as LlamaGen, and the entire training process lasts for 50 epochs. Additionally, classifier-free guidance [18] is used in the AR generation process.
Evaluation metrics.
To estimate image generation performance, we use Fréchet inception distance (FID)[17] and Inception Score (IS)[42] as the main metrics to measure the generation quality of different models. In addition, Precision and Recall[25] are also reported as secondary metrics.
4.2 Main Results
Class-conditional Image Generation. In line with LlamaGen [46], we evaluate our autoregressive generation models with parameters of  [^1] M (Hita-B), 343M(Hita-L), 775M (Hita-XL) and 1.4B (Hita-XXL) on ImageNet [8] $256\times 256$ class-conditional generation task, and compare them with the mainstream generation model families, including generative adversarial networks (GAN) [21, 43], diffusion models (Diff.) [10, 19, 41, 32], BERT-style masked-prediction models (Mask.) [5], and AR generation models [59, 12, 54, 26, 39, 46].
As shown in Table 1, our models exhibit competitive performance across all metrics compared to mainstream image generation models. Notably, our model outperforms popular diffusion models LDM like LDM-4[41] by approximately 0.7 FID and 19.6 IS with 16.6% fewer parameters. Analogously, Hita beats BERT-style models [5] in terms of FID without the requirement of complicated sampling tuning. With comparable or even fewer parameters, our method surpasses most AR generative models [59, 12, 54, 26, 39] in both FID and IS metrics. Under the same training setting, Hita surpasses LlamaGen [46] by significant FID gains and notable IS improvements. Beside, as shown in the table, extending the training duration leads to notable performance improvements of our approach over the baseline. e.g., Hita-B outperforms LlamaGen-B [46] with gains of 1.76 in FID and 56.4 in IS. Besides, our Hita-L model achieves an FID of 2.86 at 300 epochs. Beyond longer training duration, we also train an AR generation model with 2B parameters for 50 epochs. As depicted in Table. 1, our Hita-2B exhibits competitive generation performance and even surpasses LlamaGen-3B [46] with much fewer parameters. Additionally, class-conditional image generation results are also presented in Fig. 5.
4.3 Ablation Study
In the following, we first study the impact of learnable queries and semantic injection. Secondly, we perform analysis on the selection length of $k$. Next, the different foundation model for semantics injection is discussed. Later, the semantic level of holistic tokens is also studied. Further ablations are also provided in the Appendix.
Learnable queries and Semantic Injection Analysis. In Hita, it incorporates a set of learnable queries and leverages semantic injection from the pre-trained model (e.g., DINOv2 [7]) to capture holistic information. To validate their effectiveness, we conduct component analysis on the learnable queries ${Q}$ and semantic injection $\mathcal{H}$. Specifically, We first train a tokenizer with only the proposed attention modules as the baseline. Then, we gradually add learnable queries and a pre-trained foundation model (e.g., DINOv2) to the baseline. Next, we train a Hita-B AR generation model for 50 epochs to estimate the generation quality. Moreover, we conduct linear probing on the whole latent tokens after the last attention module in the baseline. As for the comparison experiments, we only perform linear probing on the holistic tokens from the last attention module. As depicted in Table. 2, simply introducing learnable queries can significantly improve image reconstruction and generation quality. The additional features from the foundation model can further enhance the reconstruction and generation quality. Besides, as indicated by the linear probing, the learnable queries can capture better semantic information compared to the baseline. With semantics injected, this semantic-level representation can be further enhanced. This illustrates that the combination of the learnable queries and semantic injection is essential for learning better holistic information, as lacking either degrades image reconstruction and generation quality.
setup image recon. usage(%) $\uparrow$ AR gen. L.P
[…]
Baseline 441 1.31 193.3 100.0 9.37 162.6 14.2
+Queries 569 1.15 191.9 30.2 100.0 6.32 187.9 28.2
+ DINOv2 1.03 198.5 100.0 100.0 5.85 212.3 36.6
Table 2: Ablation of learnable and semantic injection from pre-trained models. The learnable queries and DINOv2 [7] feature are gradually added to the baseline model. L.P. indicates linear probing on the ImageNet [8] validation set.
Figure 5: Class-conditional image generation with our proposed Hita.
Is $k\textgreater 0$ crucial? To emphasize holistic tokens and avoid its codebook collapse, Hita first builds a 1D ordered sequence with holistic tokens first and followed by the flattened patch-level tokens. Next, it is fed into a causal transformer for token fusion. Subsequently, the last $k$ holistic tokens will combined with patch-level tokens to reconstruct the image. To verify the doubt, we first configure $k$ with different values to initialize and train the holistic tokenizer. Next, we train an AR generation model – Hita-L with fixed 128 holistic tokens in total. Consequently, both image reconstruction and generation quality are estimated. As listed in Table 3, when $k$=0, the token fusion module degrades to a vanilla causal transformer. In this case, the patch-level tokens can directly influence the reconstruction of their corresponding patches, creating a shortcut that bypasses the holistic tokens. As a result, the token fusion process ignores the holistic tokens, leading to its codebook collapse, as evidenced by the 2nd row in Table. 3. Once $k\textgreater 0$, the patch-level tokens can only affect the reconstruction of other patches and must also rely on the holistic tokens. In this scenario, the problem of holistic codebook collapse is avoided. Besides, the codebook utilization rate and the reconstruction quality can also be improved. Here, we also found that $k$ equals 4 achieves the optimal reconstruction and generation quality. Thus, we configure $k$ equals 4 in our experiments, by default.
${k}$ Image recon. code $\uparrow$ AR gen.
[…]
0 1.34 187.2 0.763% 100.0% 6.80 145.4
1 1.13 192.4 100.0% 100.0% 3.78 257.1
2 1.08 195.1 100.0% 100.0% 4.14 256.8
4 1.03 198.5 100.0% 100.0% 3.75 268.8
Table 3: Ablation study of the selection length k of holistic tokens.
Pre-trained models for semantic injection. When generating holistic features, a pre-trained foundation model (e.g., DINOv2 [7]) is adopted to inject semantic information. However, the impact of using different pre-trained models to inject semantics on image reconstruction and generation quality has not yet been explored. Therefore, here we study injecting semantic-aware features from different pre-trained models, including the vision language model (e.g., CLIP [36]) or vanilla vision backbone (e.g., ResNet [14]). Here, we initialize and train a holistic tokenizer with different pre-trained foundation models with default settings. Then we train Hita-L for default 50 epochs. As shown in Table 4 , injecting DINOv2’s semantic features yields better results. Therefore, we utilize DINOv2 to perform semantic injection as default.
model image recon. AR gen.
[…]
None 1.15 191.9 6.32 187.9
CLIP [36] 1.24 191.9 6.42 186.3
ResNet [14] 1.07 195.3 6.81 191.9
DINOv2 [7] 1.03 198.5 5.85 212.3
Table 4: Comparison of different pre-trained models leveraged to perform semantic injection with advance discriminator.
What is the semantic-level of latent tokens? To analyze the semantic-level of the learned holistic tokens, we conduct linear probing according to the setting of MAE [15] on ImageNet [8] validation set. Specifically, the holistic tokens from the final causal transformer are used. Here we focus on validation instead of engaging in a comprehensive comparison with all alternative approaches.
Compared to the vanilla tokenizer VQGAN [46], we notice that Hita achieves a better accuracy, suggesting it captures semantic information. Nevertheless, it lags behind DINOv2  [7]. Because the semantic features learned by Hita are different from DINOv2’s. As shown in Fig. 2a, Hita’s primary goal is image reconstruction, focusing more on learning relatively low-level global features like texture, material, color, and other features from pixels. Nevertheless, a pre-trained model – DINOv2 integrated in Hita can inject necessary semantic-aware features to help reconstruction. Conversely, DINOv2 tends towards understanding image content, with self-supervised ViT features containing explicit information about the semantic segmentation of an image, capturing more general and high-level features.
VQGAN [46] Hita DINOv2 [7]
Acc.@Top-1 6.9 36.6 86.7
Table 5: Linear probing accuracy comparison on ImageNet [8].
5 Conclusion
We introduce Hita, a novel holistic-to-local tokenizer for AR image generation. Beyond image reconstruction, it emerges with new features such as zero-shot style transfer and zeros-shot image in-painting. Extensive experiments conducted on class-conditional image generation tasks illustrate it can accelerate converge speed during training and significantly improve the synthesis quality. Additionally, in-depth analysis also consolidates the effectiveness of the designed holistic tokenizer.
6 Acknowlegement
This work has been supported by the National Key R&D Program of China (Grant No. 2022YFB3608300), Hong Kong Research Grant Council - Early Career Scheme (Grant No. 27209621), General Research Fund Scheme (Grant No. 17202422, 17212923, 17215025) Theme-based Research (Grant No. T45-701/22-R) and Shenzhen Science and Technology Innovation Commission (SGDX20220530111405040). Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust. We are deeply grateful to Lufan Ma for the contribution in polishing up this paper.
Appendix A Comparison with other image tokenizers.
In this subsection, we first compare Hita with the other vanilla tokenizers. Next, we depict the differences between VAR [47] in detail. Later, further discussion on token fusion modules are presented.
A.1 Comparison with vanilla image tokenizers
In Table. 6, we compare with the prevalent image tokenizer, including VQGAN [12], MaskGiT [5], ViT-VQGAN [54] and TiTok [59]. The image reconstruction quality is measured by rFID [17] and rIS [42] metrics, which are evaluated on 256$\times$256 ImageNet [8] 50k validation benchmark.
Following VQGAN [12], we adopted ${\ell}_{2}$-normalization into codebook vectors, low codebook vector dimension, and a codebook size of 16,384. Compared with other counterparts like VQGAN [12] and ViT-VQGAN [54], the proposed tokenizer represents an image with fewer tokens (569 v.s. 1024), while achieving a better reconstruction quality with 100% utilization for both holistic and patch-level codebooks. Additionally, we observed that our approach achieves a better rIS 198.5 compared with the VQGAN proposed in LlamaGen [46]. rIS quantifies the KL-divergence between the original label distribution and the logit distribution of reconstructed images after softmax normalization [42]. In other words, rIS measures the semantic consistency between the reconstructed images and the original ones. A higher rIS confirms that our holistic tokenizer is more effective at preserving the semantic consistency of the reconstructed images.
Approach f setup img recon. usage(%$\uparrow$
[…]
TiTok [59] 8,192 64 256 1.05 191.5 100.0
$\text{VQGAN}^{\text{oim.}}$ [12] 8 256 4 1024 1.44
VQGAN [12] 8192 256 1.49
ViT-VQGAN [54] 8192 32 1.28 192.3 95.0
$\text{VQGAN}^{\text{oim.}}$ [12] 16384 4 1.19
VQGAN [12] 16 1024 256 256 7.94
MaskGiT [5] 2.28
Var [47] 16 4096 32 680 0.92 196.0 100.0
RQ-VAE [26] 32 16384 256 1024 1.83
VQGAN [12] 16 16384 256 256 4.98
VQGAN [46] 8 441 1.21 189.1 99.2
VQGAN [46] 8 576 0.95 197.3 99.7
Hita 8/12 569 1.03 198.5 100.0 100.0
Table 6: Comparison with other image tokenizers. ${}^{\text{oim.}}$ indicates the tokenizer is trained on OpenImages [24]. $\mathcal{Q}_{H}$/$\mathcal{Q}_{P}$ denotes the codebook usage in holistic and patch-level quantizer, respectively.
A.2 Comparison with VAR
As discussed in Sec.2, VAR’s multi-scale tokens can also be considered as a combination of semantic tokens and patch tokens. However, as depicted in Table. 7 we find removing the initial coarse-scale tokens seldom effect its reconstruction. Meanwhile, linear probing conducted on the cumulative coarse-scale tokens reveals poor semantic information.
$n$ 0 1 2 3 4 5 6 7 8
rFID 1.31 1.31 1.32 1.31 1.41 1.78 3.22 10.42 92.3
rIS 198.6 199.3 198.8 198.9 196.4 190.2 171.8 119.8 16.4
Acc 2.2 4.9 7.2 8.3 9.2 9.5 9.8 10.1 10.3
Table 7: Analysis on VAR’s multi-scale tokens. Acc indicates top-1 accuracy for linear probing estimation on the ImageNet [8].
Appendix B Further explanation on token fusion module
Figure 6: Token fusion module comparison. (a).Token fusion module composed of vanilla attention layers. (b). Our designed token fusion layers.
As shown in Fig. 6, If a vanilla transformer is used to construct the token fusion module, due to the existence of the skip connection and the patch-level tokens contain enough information for image reconstruction, the patch-level tokens can directly flow through the skip connections into the decoder to reconstruct the image. Thus, the token fusion module can learn a trivial solution, which overlooks the holistic tokens and leads to holistic codebook collapse (Fig. 6a). Once the last $k$ holistic tokens participate in the image reconstruction, the information of the patch tokens flowing through the skip connection is incomplete. It needs to interact with the holistic tokens to obtain complete information for image reconstruction(Fig. 6b). Through this simple operation, we can not only emphasize the holistic tokens but also avoid its codebook collapse. To better align the token sequence with the nature of the AR generation model, here we adopt causal attention to build the token fusion module.
Appendix C More ablation study.
In this subsection, we elaborate more ablation studies on the design of our approach. Next, we conduct a quantitative analysis of the zero-shot in-painting performance of Hita.
Figure 7: Quantitative zero-shot in-painting analysis with Hita-L conducted on ImageNet [8] evaluation benchmark.
C.1 AR generation with resolution of $256\times 256$
Given that Hita ensures a fair comparison with other approaches by controlling the number of tokens, here we directly train the image tokenizer and the AR generation model on $256\times 256$ images, enabling them to reconstruct and generate $256\times 256$ images, respectively, which is also in line with common practice. We initialize and train the image tokenizer for 40 epochs, and train both Hita-B and Hita-L for 50 epochs as default. As depicted in Table. 8, Hita can not only achieve a better reconstruction performance, but also improve the generation quantity compare to LlamaGen [46].
Approach Image recon. code $\uparrow$ AR gen.
[…]
LlamaGen-B 2.22 169.8 95.2% 7.22 178.3
LlamaGen-L 4.21 200.0
Hita-B 1.40 186.6 100% 100% 6.58 210.2
Hita-L 4.04 242.2
Table 8: Hita performs image reconstruction and AR generation with image resolution of $256\times 256$.
C.2 Other token fusion variants
As depicted in Sec 3.2.3, for simplicity, we choose the last $k$ holistic tokens combined with the patch-level tokens to reconstruct the image. Here, we refer to MAE [15] and Titok [59] and design 2 different variants to generate features for the first $k$ patch tokens: 1). Partial: A mask token combined with the positional embedding of the first $k$ patch tokens are used to generate their feature. In this scenario, the holistic tokens can be treated as condition; 2). Full: The patch tokens are completely removed, which is consistent with TiTok [59]. Then a mask token along with the positional embeddings is used to generate the features for all patch tokens. To achieve this, we initialize and train 2 different holistic tokenizers for 40 epochs, and then train 2 AR generation models – Hita-B and Hita-L based on those tokenizers with default training settings for 50 epochs. As listed in Table 9, the token fusion strategy proposed in Hita shows a better performance in both reconstruction and generation compared to the other variants. Thus, we choose to use the last ${k}$ holistic tokens combined with the patch-level tokens to reconstruct the image, by default.
Variants image recon. code $\uparrow$ AR gen.
[…]
Partial 1.05 198.2 100.0% 100.0% B 6.59 209.8
L 3.96 243.1
Full 2.07 170.3 100.0% B 11.64 172.1
L 6.75 219.7
Hita 1.03 198.5 100.0% 100.0% B 5.85 212.3
L 3.75 262.1
Table 9: Image reconstruction and AR generation. with other token fusion strategies.
C.3 Attention Modules Study.
[…]
As depicted in Table. 10, it can be observed when only $\mathcal{E}_{\text{trans}}$ is directly discarded from the tokenizer, the quality of image reconstruction and generation slightly drops. This is because the subsequent causal transformer $\mathcal{E}_{\text{causal}}$ can simultaneously achieve the requirement of holistic feature capture, semantic-aware feature injection, and feature space alignment. Similarly, only removing $\mathcal{E}_{\text{causal}}$ leads to a slight degradation in image reconstruction and generation, indicating that incorporating causal attention $\mathcal{E}_{\text{trans}}$ within the tokenizer helps in learning a latent space that better aligns with the causal nature of AR models. With all the modules integrated, we achieve the best performance in terms of reconstruction and generation.
[…]
1.21 189.1 99.7% 7.95 166.9
1.15 196.0 100.0% 100.0% 6.01 196.3
1.20 195.3 100.0% 100.0% 6.12 188.9
1.03 198.5 100.0% 100.0% 5.85 212.3
Table 10: Attention modules study in Hita. ‘-’ indicates the module is discarded from the tokenizer’s architecture.
C.4 Quantitative study on the in-painting quality.
Beyond the demonstration that the holistic tokenizer can help AR generation model maintain a better semantic consistency (see Fig. 9), we conduct a quantitative analysis of this. Similar to the experimental setup described in the manuscript, we only retain a certain fraction of upper part image, e.g. $25\%$, $50\%$, etc. then utilize a tokenizer to discretize it into visual tokens. These tokens are fed as a prefix sequence prompt to a pre-trained AR generation model, which is required to complete the lower part of the image. Here, we take the AR generation models – Hita-L and LlamaGen-L [46] for a fair comparison.
Evaluation metrics. To quantitatively estimate the semantic consistency of the generated images, we adopt top-1 accuracy and CLIP-score as our metrics, along with the generation metrics FID and IS. The top-1 accuracy is derived from image classification tasks, where we utilize a pre-trained ResNet-101 [14] on ImageNet [8] to classify the completed images. CLIP-score measures the similarity between the original image and the completed one. Specifically, we feed both original and in-painted images into the CLIP [36] model to extract image features and compute their cosine distance averaged across all samples. Larger top-1 accuracy and CLIP-score indicate a better semantic consistency is maintained in an AR generation model.
Observation and discussion. As shown in Fig. 7 and Fig. 9, the AR generation model trained with vanilla VQGAN [12] encounters difficulties in producing an image that maintains semantic coherence. In contrast, our approach can effectively produce visually consistent content to complete the given image part, maintaining a better overall coherence even with a significantly truncated prefix. Specifically, as observed in Fig. 7.a and Fig. 7.b, our method achieves higher classification accuracy and CLIP-score under various settings, with relatively stable fluctuations in all metrics across different configurations. In contrast, models trained with vanilla VQGAN exhibit more obvious performance variations, especially when only a small portion of the image is provided.
Beyond semantic coherency measurement, we also estimate the completed image quality with generation evaluation metrics, FID and IS. As shown in Fig. 7.c and Fig. 7.d, compared to the model trained with vanilla VQGAN[12], our approach achieves a better FID and IS, which also illustrates the holistic tokenizer can help the AR generation model generate the lower half of images robustly, maintaining strong semantic consistency.
Appendix D Limitation
Currently, Hita is trained on a limited dataset with vanilla optimization techniques, better performance could be achieved with more data and advanced learning objectives. Additionally, Hita can be seamlessly extended to perform text-conditional image generation, which is an ongoing direction of our research.
Appendix E More Visual Samples
In this section, we present more visualization samples including zero-shot style transfer (see Fig. 8), image in-painting (see Fig. 9), and class-conditional image generation (see Fig. 10). For optimal clarity, please zoom in.
Figure 8: Some zero-shot style-transfer samples by Hita’s holistic tokenizer. Best viewed with zoom-in.
Figure 9: Zero-shot in-painting examples by Hita’s AR generation model. Compared with the baseline. Best viewed with zoom-in.
Figure 10: Visualization of class-conditional samples generated by Hita. Best viewed with zoom-in.

### v4

@@
 Abstract
-The vanilla autoregressive image generation model generates visual tokens in a step-by-step fashion, which limits the ability to capture holistic relationships among token sequences.
-Moreover, most visual tokenizers map local image patches into latent tokens, leading to limited global information.
+Vanilla autoregressive image generation models generate visual tokens step-by-step, limiting their ability to capture holistic relationships among token sequences.
+Moreover, because most visual tokenizers map local image patches into latent tokens, global information is limited.
 To address this, we introduce Hita, a novel image tokenizer for autoregressive (AR) image generation.
 It introduces a holistic-to-local tokenization scheme with learnable holistic queries and local patch tokens.
-Besides, Hita incorporates two key strategies for improved alignment with the AR generation process: 1) it arranges a sequential structure with holistic tokens at the beginning followed by patch-level tokens while using causal attention to maintain awareness of previous tokens; and 2) before feeding the de-quantized tokens into the decoder, Hita adopts a lightweight fusion module to control information flow to prioritize holistic tokens.
+Hita incorporates two key strategies to better align with the AR generation process: 1) arranging a sequential structure with holistic tokens at the beginning, followed by patch-level tokens, and using causal attention to maintain awareness of previous tokens; and 2) adopting a lightweight fusion module before feeding the de-quantized tokens into the decoder to control information flow and prioritize holistic tokens.
 Extensive experiments show that Hita accelerates the training speed of AR generators and outperforms those trained with vanilla tokenizers, achieving 2.59 FID and 281.9 IS on the ImageNet benchmark.
-A detailed analysis of the holistic representation highlights its ability to capture global image properties such as textures, materials, and shapes.
+Detailed analysis of the holistic representation highlights its ability to capture global image properties, such as textures, materials, and shapes.
 Additionally, Hita also demonstrates effectiveness in zero-shot style transfer and image in-painting.
@@
 This challenge has also been highlighted in works such as [11, 27, 53], where bi-directional attention mechanisms are introduced to mitigate the issue, albeit with non-trivial modifications to the AR generation process.
-To address this, we introduce Hita, a novel global-to-local image tokenizer for AR image generation.
-First, we propose to utilize a set of learnable queries to capture the holistic information from the patch-level embeddings of an image.
+To address these issues, we introduce Hita, a novel global-to-local image tokenizer for AR image generation.
+First, we propose using learnable queries to capture the holistic information from the patch-level image embeddings.
 Meanwhile, a pre-trained foundation model, such as DINOv2 [7], is adopted to inject semantics into the learned queries to enhance their representation.
-Then, we utilize 2 separate codebooks to quantize the learned queries and patch-level embeddings, respectively, resulting in holistic tokens and patch-level tokens.
-Next, we fuse the de-quantized holistic and patch-level tokens together with a novel token fusion module to reconstruct the image.
-Such design can better prioritize the holistic token and improve both image reconstruction and generation quality.
+Then, we quantize the learned queries and patch-level embeddings using two separate codebooks, resulting in holistic tokens and patch-level tokens.
+Next, we fuse the de-quantized holistic and patch-level tokens using a novel token fusion module to reconstruct the image.
+This design can better prioritizes the holistic token and improving both image reconstruction and generation quality.
 Once Hita is trained, a vanilla autoregressive (AR) transformer model, e.g.
 Llama [48], can be seamlessly incorporated to generate both holistic and image patch tokens sequentially, which can be decoded into image facilitated by Hita.
-During the AR generation, the AR generator first generates holistic tokens, which encapsulate the global image information and serve as a prefix prompt.
-These holistic tokens then guide the subsequent generation of local image patch tokens, alleviating difficulties and maintaining global coherence to improve generation quality.
+During the AR generation, the AR generator first generates holistic tokens, which encapsulate global image information and serve as a prefix prompt.
+These holistic tokens then guide the subsequent generation of local image patch tokens, alleviating difficulties and maintaining global coherence for improved generation quality.
 Our in-depth ablation experiments confirm that the learned holistic tokens effectively capture global image features, such as shape, color, texture, etc, from pixels (see Fig.
 2a).
-This has been demonstrated through effortlessly enabling style transfer by combining holistic tokens from reference images with patch-level tokens from the source image.
+This is demonstrated by effortlessly enabling style transfer by combining holistic tokens from reference images with patch-level tokens from the source image.
 Moreover, without task-specific training, the model can be seamlessly applied to image inpainting by encoding the given incomplete image into latent tokens, which serve as prefix sequence prompts to the AR model, requiring it to generate visual tokens for completing the image (see Fig.
@@
 It also surpass the popular diffusion model LDM-4[41] by 0.9 FID and 27.1 IS, respectively.
-To summarize, our contribution can be listed as follow:
-  • We propose a novel global-to-local image tokenizer for AR image generation, with which an AR generation model can be integrated without any modification.
-  • The proposed image tokenizer emerges with new characteristics, such as zero-shot style stransfer and zero-shot image in-painting.
-  • Extensive experiments conducted on class-conditional image generation task demonstrates it can accelerate the converge speed during training and significantly improve the synthesis quality.
+In summary, our contributions are:
+  • We propose a novel global-to-local image tokenizer for AR image generation that allows integration of AR generation models without modification.
+  • The proposed image tokenizer emerges with new characteristics, such as zero-shot style transfer and zero-shot image in-painting.
+  • Experiments on class-conditional image generation demonstrate that it accelerates convergence during training and significantly improves synthesis quality.
 2 Related Work
@@
 TiTok [59] and VAR [47] are related efforts but differ from Hita in several aspects.
-TiTok aims to create compact 1D tokens with queries to reduce redundancy, while Hita uses queries to capture holistic properties for 1D holistic tokens and fuses them with 2D image patch tokens to reconstruct images.
-TiTok refines queries for 1D tokens tailored for reconstruction objectives, whereas Hita extracts information from pre-trained foundation model and patch features to capture global information.
-TiTok requires a complex training strategy, while Hita enables efficient single-stage training.
+TiTok creates compact 1D tokens with queries to reduce redundancy, while Hita uses queries to capture holistic properties for 1D holistic tokens and fuses them with 2D image patch tokens for image reconstruction.
+TiTok refines queries for 1D tokens tailored for reconstruction, whereas Hita extracts information from a pre-trained foundation model and patch features to capture global information.
+TiTok requires a complex training strategy, while Hita allows efficient single-stage training.
 Hita outperforms TiTok in AR generation and offers new features like zero-shot style transfer and inpainting.
 VAR [47] predicts global and local tokens indiscriminately, while Hita separates holistic and patch-level tokens, enriching holistic tokens with pre-trained foundation to form global representations and uses causal attention to align with the causal nature of AR models, which VAR lacks.
-Removing initial coarse-scale tokens seldom affects VAR’s reconstruction, indicating they are not crucial.
+Removing initial coarse-scale tokens seldom affects VAR’s reconstruction, indicating those tokens are not crucial.
 Autoregressive Image Generation.
@@
 3 Method
-In this section, we first introduce the preliminary related to image tokenizer and AR generation models.
-Then, we present the design of our holistic tokenizer step by step.
-Next, their characteristics are also discussed.
+In this section, we first introduce the preliminaries related to image tokenizers and AR generation models.
+Then, we present the design of our holistic tokenizer step-by-step.
+Next, we discuss their characteristics.
 Figure 3: The flowchart of the proposed holistic image tokenizer, which includes: Holistic feature extraction part incorporates learnable queries to obtain holistic and patch-level features from pixels.
@@
 raster-scan order.
-Then, an AR generation model predicts the sequence of image tokens autoregressively.
-To achieve this, AR models typically incorporate causal attention with prefix token sequence to predict the next, which focuses on learning the dependencies between the prefix discrete image tokens and the current one.
+Then, an AR generation model autoregressively predicts the sequence of image tokens.
+To achieve this, AR models typically incorporate causal attention, using a prefix token sequence to predict the next, focusing on learning the dependencies between the prefix and current discrete image tokens.
 Considering a sequence of discrete tokens ${x_{1:N}}$, an AR model is trained on these tokens by maximizing the cross-entropy depicted in Eq.
@@
 $\mathop{\max}\mathcal{L}^{\text{CE}}_{\theta}=\sum\limits_{i=1}^{N}\text{log}P_{\theta}({x_{i}}|{x_{j<i}},c)$ (3)
-During inference, the AR model initially predicts visual tokens by sampling tokens from the learned distribution autoregressively.
+During inference, the AR model initially predicts visual tokens autoregressively by sampling tokens from the learned distribution.
 Subsequently, the generated visual tokens are converted back into images through the decoder $\mathcal{D}(\cdot)$.
@@
 3, Hita incorporates the following key components.
-First, to capture holistic property, we introduce learnable queries that interact with image patch embeddings via attention modules.
+First, to capture holistic properties, we introduce learnable queries that interact with image patch embeddings via attention modules.
 A pre-trained foundation model (e.g., DINOv2 [7]) injects semantic-level features.
-To align Hita’s latent space with the causal nature of AR models, we use a causal transformer to integrate holistic and patch features before quantization, embedding the sequential structure into the tokenization process to learn prefix-aware latent tokens.
+To align Hita’s latent space with the causal nature of AR models, we use a causal transformer to integrate holistic and patch features before quantization, embedding the sequential structure into the tokenization process for learning prefix-aware latent tokens.
 Details are depicted in Sec.
@@
 3.2.2).
-Finally, in the token fusion and decoding procedure after de-quantization, Hita first integrates a fusion module with casual transformers to process these tokens in order and prioritize holistic tokens.
+Finally, after dequantization and during token fusion and decoding procedure, Hita first uses a fusion module with causal transformers to process these tokens in order and prioritize holistic tokens.
 Then they are fed into the decoder for image reconstruction (see Sec.
@@
 We build Hita based on VQGAN [12].
-Specifically, we adopt its encoder $\mathcal{E}(\cdot)$ to encode input image $I$ into local image patch embeddings, and further incorporate a set of learnable queries ${Q}$ to gather holistic information from the whole input.
+Specifically, we adopt its encoder $\mathcal{E}(\cdot)$ to encode input image $I$ into local patch embeddings, and further incorporate a set of learnable queries ${Q}$ to gather holistic information from the whole input.
 To enhance the holistic representation, a pre-trained foundation model $\mathcal{H}(\cdot)$ like DINOv2 [7] is leveraged to inject rich semantic-level features.
 To achieve this, we first concatenate the learnable queries and the flattened patch-level embeddings as well as foundation model features into sequence and feed them into a simple transformer $\mathcal{E}_{\text{{trans}}}(\cdot)$ is utilized, where holistic queries ${Q}$ interact with all patch-level image embeddings and the injected semantic-level features.
-Considering the AR model models the probability of observing the current token based on its predecessors, a well-aligned latent space with the causal nature is encouraged.
-Therefore, the extracted holistic feature and patch-level feature are concatenated and reorganized in an ordered sequence that holistic queries come first followed by the patch-level image patch feature obtained in raster-scan order.
+Since the AR model models the probability of observing the current token based on its predecessors, a latent space well-aligned with the causal nature is encouraged.
+Therefore, the extracted holistic and patch-level features are concatenated and reorganized in an ordered sequence, with holistic queries come first followed by the patch-level features obtained in raster-scan order.
 This sequence is then fed into a transformer with global features serving as prefix prior.
@@
 Moreover, the pre-trained feature is discarded after $\mathcal{E}_{\text{trans}}$.
-Figure 4: The illustration of token fusion and decoding procedure
+Figure 4: Token fusion and decoding procedure
 3.2.2 Separate Vector Quantization
@@
 Then, the holistic tokens $\hat{Q}$ and patch-level $\hat{Z}$ tokens could be obtained.
-Considering the codebook significantly impacts the performance of the image tokenizer.
+The the codebook significantly impacts the performance of the image tokenizer.
 Following [54, 46], we apply ${\ell}_{2}$-normalization to the codebook vectors, opting for a lower vector dimension and a larger codebook size to improve the reconstruction quality and codebook utilization.
 3.2.3 Token Fusion and Decoding Procedure
-After quantization, the holistic tokens and patch-level tokens could be fed into a vanilla transformer to perform token fusion and then fed into the decoder to reconstruct the image.
+After quantization, the holistic tokens and patch-level tokens can be fed into a standard transformer for token fusion and then fed into the decoder to reconstruct the image.
 However, such a design incurs codebook collapse in the holistic quantizer (3th row in Table.
@@
 1a in Appendix).
-As a result, the token fusion module learned a trivial solution, which overlooks the holistic tokens and leads to holistic codebook collapse.
+As a result, the token fusion module learns a trivial solution that overlooks the holistic tokens and leads to holistic codebook collapse.
 […]
@@
 In this case, we set both $\alpha$ and $\lambda$ to 1.
-3.3 Autoregresive Image Generation
+3.3 Autoreggresive Image Generation
 Once the tokenizer is trained, a standard autoregressive (AR) transformer model, e.g.
@@
 Starting from the class/text embedding $c$, the AR model first generates holistic tokens as prefix prompts, then guides the subsequent generation of patch token sequence in the way of next-token prediction.
-Once all the visual tokens are generated, they will be tackled by the token fusion and decoding procedure to be converted into pixels.
-Additionally, in AR models, each visual token is added with positional embedding.
+Once all the visual tokens are generated, they undergo the token fusion and decoding procedure to be converted into pixels.
+Additionally, in AR models, each visual token receives a positional embedding.
 Since holistic tokens form a 1D sequence, we add learnable positional embeddings for them.
@@
 2a, the holistic tokens can capture global information from an image, such as texture, color, materials, shapes, and other features from the pixels.
-For instance, in the case of the lizard image, Hita successfully transfers the material properties of the ship from the reference image in thumbnail to the input image, creating a realistic mechanical lizard.
+For instance, in the case of the lizard image, Hita successfully transfers the material properties of the ship from the reference image to the input image, creating a realistic mechanical lizard.
 More samples are shown in Appendix.
@@
 This setup prevents local information leakage during in-painting.
-We first tokenize the images into holistic and patch tokens, then feed them to the AR model to generate additional tokens, conditioned on class embeddings in AR model.
+We first tokenize the incompleted images into holistic and patch tokens, then feed them to the AR model, conditioned on class embeddings, to generate additional tokens.
 Hita-XXL is used without modifications to the model architecture or parameter tuning.
@@
 For example, in the case of the fish image, Hita correctly generates a complete fish, while LlamaGen [46] mistakenly generates a “fish-bird”.
-These results confirm that our holistic tokenizer is semantic-aware and has a stronger ability to generate semantically and holistically coherent examples.
+These results confirm that our holistic tokenizer is semantic-aware and has a stronger ability to better generate semantically and holistically coherent examples.
 Furthermore, the ablation study of linear probing on holistic tokens is conducted in Sec.
@@
 To ensure a fair comparison with LlamaGen [46], we train the tokenizer on $336\times 336$ images, which yields a comparable number of discrete tokens to it.
-We also adopt the training settings for tokenizer training from LlamaGen.
+We also adopt the tokenizer training settings from LlamaGen.
 Additionally, during evaluation, the test images are resized to $256\times 256$, consistent with the evaluation procedure in LlamaGen.
@@
 Following the AR generation procedure in LlamaGen [46], the AR models first generate images of the same size as the image tokenizer and then resize them to $256\times 256$ for evaluation.
-Unless otherwise specified, all models are trained with the same settings as LlamaGen, and the entire training process lasts for 50 epochs.
+Unless otherwise specified, all models are trained with the same settings as LlamaGen, with the entire training process lasting 50 epochs.
 Additionally, classifier-free guidance [18] is used in the AR generation process.
@@
 4.3 Ablation Study
-In the following, we first study the impact of learnable queries and semantic injection.
-Secondly, we perform analysis on the selection length of $k$.
-Next, the different foundation model for semantics injection is discussed.
-Later, the semantic level of holistic tokens is also studied.
-Further ablations are also provided in the Appendix.
+First, we study the impact of learnable queries and semantic injection.
+Second, we analyze the selection length $k$.
+Next, we discuss different foundation models for semantic injection.
+Later, we study the semantic level of holistic tokens.
+We provide further ablation in the Appendix.
 Learnable queries and Semantic Injection Analysis.
-In Hita, it incorporates a set of learnable queries and leverages semantic injection from the pre-trained model (e.g., DINOv2 [7]) to capture holistic information.
-To validate their effectiveness, we conduct component analysis on the learnable queries ${Q}$ and semantic injection $\mathcal{H}$.
-Specifically, We first train a tokenizer with only the proposed attention modules as the baseline.
-Then, we gradually add learnable queries and a pre-trained foundation model (e.g., DINOv2) to the baseline.
+In Hita, it incorporates a set of learnable queries and leverages semantic injection from a pre-trained model (e.g., DINOv2 [7]) to capture holistic properties.
+To validate their effectiveness, we conduct analysis on the learnable queries ${Q}$ and semantic injection $\mathcal{H}$.
+Specifically, We first train a tokenizer with only the attention modules as a baseline.
+Then, we gradually add learnable queries and a pre-trained foundation model (e.g., DINOv2) to this baseline.
 Next, we train a Hita-B AR generation model for 50 epochs to estimate the generation quality.
@@
 The additional features from the foundation model can further enhance the reconstruction and generation quality.
-Besides, as indicated by the linear probing, the learnable queries can capture better semantic information compared to the baseline.
+Besides, as indicated by the linear probing, the learnable queries can capture better semantic information than the baseline.
 With semantics injected, this semantic-level representation can be further enhanced.
-This illustrates that the combination of the learnable queries and semantic injection is essential for learning better holistic information, as lacking either degrades image reconstruction and generation quality.
+This illustrates that combining the learnable queries and semantic injection is essential for learning better holistic information, as lacking either degrades image reconstruction and generation quality.
 setup image recon. usage(%) $\uparrow$ AR gen. L.P
@@
 Figure 5: Class-conditional image generation with our proposed Hita.
-Is $k\textgreater 0$ crucial?
+Is $k>0$ crucial?
 To emphasize holistic tokens and avoid its codebook collapse, Hita first builds a 1D ordered sequence with holistic tokens first and followed by the flattened patch-level tokens.
 Next, it is fed into a causal transformer for token fusion.
-Subsequently, the last $k$ holistic tokens will combined with patch-level tokens to reconstruct the image.
-To verify the doubt, we first configure $k$ with different values to initialize and train the holistic tokenizer.
+Subsequently, the last $k$ holistic tokens are combined with patch-level tokens to reconstruct the image.
+To verify the doubt, we first initialize and train the holistic tokenizer, configuring $k$ with different values.
 Next, we train an AR generation model – Hita-L with fixed 128 holistic tokens in total.
@@
 3.
-Once $k\textgreater 0$, the patch-level tokens can only affect the reconstruction of other patches and must also rely on the holistic tokens.
+Once $k>0$, the patch-level tokens can only affect the reconstruction of other patches and must also rely on the holistic tokens.
 In this scenario, the problem of holistic codebook collapse is avoided.
-Besides, the codebook utilization rate and the reconstruction quality can also be improved.
+This also improves the codebook utilization and the reconstruction quality.
 Here, we also found that $k$ equals 4 achieves the optimal reconstruction and generation quality.
-Thus, we configure $k$ equals 4 in our experiments, by default.
+Thus, we set $k$ to 4 by default in our experiments.
 ${k}$ Image recon. code $\uparrow$ AR gen.
@@
 When generating holistic features, a pre-trained foundation model (e.g., DINOv2 [7]) is adopted to inject semantic information.
-However, the impact of using different pre-trained models to inject semantics on image reconstruction and generation quality has not yet been explored.
-Therefore, here we study injecting semantic-aware features from different pre-trained models, including the vision language model (e.g., CLIP [36]) or vanilla vision backbone (e.g., ResNet [14]).
-Here, we initialize and train a holistic tokenizer with different pre-trained foundation models with default settings.
+However, the impact of using different pre-trained models for semantic injection on image reconstruction and generation remains unexplored.
+Here we study injecting semantic-aware features from different pre-trained models, including the vision language model (e.g., CLIP [36]) or vanilla vision backbone (e.g., ResNet [14]).
+We initialize and train a holistic tokenizer with different pre-trained foundation models using the default settings.
 Then we train Hita-L for default 50 epochs.
@@
 DINOv2 [7] 1.03 198.5 5.85 212.3
-Table 4: Comparison of different pre-trained models leveraged to perform semantic injection with advance discriminator.
+Table 4: Comparison of semantic injection performance with different pre-trained models.
 What is the semantic-level of latent tokens?
@@
 Specifically, the holistic tokens from the final causal transformer are used.
-Here we focus on validation instead of engaging in a comprehensive comparison with all alternative approaches.
+Here we focus on estimation rather than engaging in a comprehensive comparison with all alternative approaches.
 Compared to the vanilla tokenizer VQGAN [46], we notice that Hita achieves a better accuracy, suggesting it captures semantic information.
@@
 2a, Hita’s primary goal is image reconstruction, focusing more on learning relatively low-level global features like texture, material, color, and other features from pixels.
-Nevertheless, a pre-trained model – DINOv2 integrated in Hita can inject necessary semantic-aware features to help reconstruction.
+Nevertheless, DINOv2, a pre-trained model integrated in Hita, injects necessary semantic-aware features to aid reconstruction.
 Conversely, DINOv2 tends towards understanding image content, with self-supervised ViT features containing explicit information about the semantic segmentation of an image, capturing more general and high-level features.
@@
 5 Conclusion
-We introduce Hita, a novel holistic-to-local tokenizer for AR image generation.
-Beyond image reconstruction, it emerges with new features such as zero-shot style transfer and zeros-shot image in-painting.
-Extensive experiments conducted on class-conditional image generation tasks illustrate it can accelerate converge speed during training and significantly improve the synthesis quality.
-Additionally, in-depth analysis also consolidates the effectiveness of the designed holistic tokenizer.
-6 Acknowlegement
+Autoregressive (AR) models struggle with image generation because their unidirectional attention mechanisms restrict their ability to capture global context, resulting in lower generation quality.
+Hita addresses this by introducing a holistic-to-local tokenization scheme that uses learnable holistic tokens and local patch tokens to incoporate global information.
+Beyond image tokenization, Hita exhibits new features such as zero-shot style transfer and zeros-shot image in-painting.
+Experiments conducted on class-conditional image generation tasks illustrate that it accelerates convergence speed during training and significantly improves the synthesis quality.
+Besides, in-depth analysis also consolidates its effectiveness.
+Acknowledgments
 This work has been supported by the National Key R&D Program of China (Grant No.
@@
 Part of the described research work is conducted in the JC STEM Lab of Robotics for Soft Materials funded by The Hong Kong Jockey Club Charities Trust.
-We are deeply grateful to Lufan Ma for the contribution in polishing up this paper.
+We are deeply grateful to Lufan Ma for the contribution in editing this paper.
 Appendix A Comparison with other image tokenizers.
-In this subsection, we first compare Hita with the other vanilla tokenizers.
-Next, we depict the differences between VAR [47] in detail.
-Later, further discussion on token fusion modules are presented.
+In this subsection, we first compare Hita with the other vanilla tokenizers. Next, we depict the differences between VAR [47] in detail. Next, we present further discussion on token fusion modules.
 A.1 Comparison with vanilla image tokenizers
@@
 Following VQGAN [12], we adopted ${\ell}_{2}$-normalization into codebook vectors, low codebook vector dimension, and a codebook size of 16,384.
-Compared with other counterparts like VQGAN [12] and ViT-VQGAN [54], the proposed tokenizer represents an image with fewer tokens (569 v.s.
+Compared with other counterparts like VQGAN [12] and ViT-VQGAN [54], the proposed tokenizer represents an image with fewer tokens (569 vs.
 1024), while achieving a better reconstruction quality with 100% utilization for both holistic and patch-level codebooks.
@@
 Table 6: Comparison with other image tokenizers.
-${}^{\text{oim.}}$ indicates the tokenizer is trained on OpenImages [24].
-$\mathcal{Q}_{H}$/$\mathcal{Q}_{P}$ denotes the codebook usage in holistic and patch-level quantizer, respectively.
+${}^{\text{oim.}}$ indicates training on OpenImages [24].
+$\mathcal{Q}_{H}$/$\mathcal{Q}_{P}$ denote the codebook usage in holistic and patch-level quantizers, respectively.
 A.2 Comparison with VAR
@@
 However, as depicted in Table.
-7 we find removing the initial coarse-scale tokens seldom effect its reconstruction.
+7 we find removing the initial coarse-scale tokens seldom effects its reconstruction.
 Meanwhile, linear probing conducted on the cumulative coarse-scale tokens reveals poor semantic information.
@@
 Acc 2.2 4.9 7.2 8.3 9.2 9.5 9.8 10.1 10.3
-Table 7: Analysis on VAR’s multi-scale tokens. Acc indicates top-1 accuracy for linear probing estimation on the ImageNet [8].
-Appendix B Further explanation on token fusion module
+Table 7: Analysis of VAR’s multi-scale tokens. Acc indicates top-1 accuracy for linear probing estimation on the ImageNet [8].
+Appendix B Explanation on Token Fusion Module
 Figure 6: Token fusion module comparison. (a).Token fusion module composed of vanilla attention layers. (b). Our designed token fusion layers.
@@
 6b).
-Through this simple operation, we can not only emphasize the holistic tokens but also avoid its codebook collapse.
+This simple operation emphasizes the holistic tokens and avoids codebook collapse.
 To better align the token sequence with the nature of the AR generation model, here we adopt causal attention to build the token fusion module.
-Appendix C More ablation study.
-In this subsection, we elaborate more ablation studies on the design of our approach. Next, we conduct a quantitative analysis of the zero-shot in-painting performance of Hita.
+Appendix C More Ablation Studies.
+We elaborate on further ablation studies on the design of our approach. Next, we quantitatively analyze Hita’s zero-shot inpainting performance.
 Figure 7: Quantitative zero-shot in-painting analysis with Hita-L conducted on ImageNet [8] evaluation benchmark.
@@
 Full: The patch tokens are completely removed, which is consistent with TiTok [59].
-Then a mask token along with the positional embeddings is used to generate the features for all patch tokens.
+Then, a mask token, along with the positional embeddings, generates the features for all patch tokens.
 To achieve this, we initialize and train 2 different holistic tokenizers for 40 epochs, and then train 2 AR generation models – Hita-B and Hita-L based on those tokenizers with default training settings for 50 epochs.
@@
 L 3.75 262.1
-Table 9: Image reconstruction and AR generation. with other token fusion strategies.
+Table 9: Image reconstruction and AR generation. with different token fusion strategies.
 C.3 Attention Modules Study.
@@
 1.03 198.5 100.0% 100.0% 5.85 212.3
-Table 10: Attention modules study in Hita. ‘-’ indicates the module is discarded from the tokenizer’s architecture.
-C.4 Quantitative study on the in-painting quality.
+Table 10: Attention modules analysis in Hita. ‘-’ indicates the module was removed from the tokenizer’s architecture.
+C.4 Quantitative Study of In-painting Quality.
 Beyond the demonstration that the holistic tokenizer can help AR generation model maintain a better semantic consistency (see Fig.
@@
 The top-1 accuracy is derived from image classification tasks, where we utilize a pre-trained ResNet-101 [14] on ImageNet [8] to classify the completed images.
-CLIP-score measures the similarity between the original image and the completed one.
+CLIP-score measures the similarity between the original and the completed images.
 Specifically, we feed both original and in-painted images into the CLIP [36] model to extract image features and compute their cosine distance averaged across all samples.
-Larger top-1 accuracy and CLIP-score indicate a better semantic consistency is maintained in an AR generation model.
+Higher top-1 accuracy and CLIP-scores indicate a better semantic consistency is maintained in an AR generation model.
 Observation and discussion.
@@
 9, the AR generation model trained with vanilla VQGAN [12] encounters difficulties in producing an image that maintains semantic coherence.
-In contrast, our approach can effectively produce visually consistent content to complete the given image part, maintaining a better overall coherence even with a significantly truncated prefix.
+In contrast, our approach effectively produces visually consistent content to complete the given image part, maintaining a better overall coherence even with a significantly truncated prefix.
 Specifically, as observed in Fig.
@@
 In contrast, models trained with vanilla VQGAN exhibit more obvious performance variations, especially when only a small portion of the image is provided.
-Beyond semantic coherency measurement, we also estimate the completed image quality with generation evaluation metrics, FID and IS.
+Beyond measuring semantic coherence, we also estimate the completed image quality using the generation evaluation metrics, FID and IS.
 As shown in Fig.
@@
 7.d, compared to the model trained with vanilla VQGAN[12], our approach achieves a better FID and IS, which also illustrates the holistic tokenizer can help the AR generation model generate the lower half of images robustly, maintaining strong semantic consistency.
-Appendix D Limitation
-Currently, Hita is trained on a limited dataset with vanilla optimization techniques, better performance could be achieved with more data and advanced learning objectives.
+Appendix D Limitations
+Currently, Hita is trained on a limited dataset using basic optimization techniques.
+Better performance could be achieved with more data and advanced learning objectives.
 Additionally, Hita can be seamlessly extended to perform text-conditional image generation, which is an ongoing direction of our research.
-Appendix E More Visual Samples
+Appendix E More Visual Examples
 In this section, we present more visualization samples including zero-shot style transfer (see Fig.